OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 + 4x − 1 in vertex form

OpenStudy (anonymous):

Group the terms like this \[f(x)=(x^2+4x)-1\] Take the coefficient of x, divide it by 2 and square it. What's \(\left( \frac{ 4 }{ 2 } \right)^2\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Right, so add 4 inside the parentheses. This means you also have to subtract 4 outside, so the equation doesn't change. \[f(x)=(x^2+4x+4)-1-4\]

OpenStudy (anonymous):

Now factor the part in parentheses, and combine like terms on the outside

OpenStudy (anonymous):

so would it be f(x)=(x^2+4x+4)-5

OpenStudy (anonymous):

yes, but you also have to factor x² + 4x + 4

OpenStudy (anonymous):

oh I got it. But can you help me with a few others?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Using the completing-the-square method, find the vertex of the function f(x) = –2x2 + 12x + 5 and indicate whether it is a minimum or a maximum and at what point.

OpenStudy (anonymous):

\(f(x)=-2x^2+12x+5\) Group it again \(f(x)=(-2x^2+12x)+5\) This time, because you have -2 in front of x², you have to factor it out. Can you do that?

OpenStudy (anonymous):

would it be f(x)=-2(x^2+12x)+5

OpenStudy (anonymous):

no, you have to divide 12 by -2 as well because it's also in parentheses

OpenStudy (anonymous):

so it would be 6?

OpenStudy (anonymous):

12/(-2)= -6 So you have \[f(x)=-2(x^2-6x)+5\] Now take the coefficient of x, that's -6. Divide it by 2, the square the result like we did above

OpenStudy (anonymous):

okay so would it be minimum or maximum?

OpenStudy (anonymous):

the number in front of x² is negative, so it's a maximum

OpenStudy (anonymous):

okay so whats the point? (-3,5)?

OpenStudy (anonymous):

You have to complete the square to find the point

OpenStudy (anonymous):

If you don't want to do it that way, use \[x=-\frac{ b }{ 2a }\]

OpenStudy (anonymous):

okay I just reset my exam so now I have different questions

OpenStudy (anonymous):

ughhhh I need help